OpenStudy (anonymous):

Given ƒ(x) = 7x + 1 and g(x) = x2, find (g ○ ƒ)(x).

OpenStudy (anonymous):

\[f\circ g(x)=f(g(x))\] always work from the inside out, it is easiest

OpenStudy (anonymous):

\[g\circ f(x)=g(f(x))=g(7x+1)\] is a start

OpenStudy (anonymous):

then since \[g(\heartsuit)=\heartsuit^2\] you have \[g(7x+1)=(7x+1)^2\]

OpenStudy (anonymous):

you can square if you like, just make sure to do it careully

OpenStudy (anonymous):

\[(7x+1)^2=(7x+1)(7x+1)\] etc